Part I Suppose that in the market for paper, demand is P=100 - Q. The marginal private cost of producing paper is 10+ Q. Howe10+Q+ 18 (atter tox) 10+Q (before tax) 0 6 45 100 Part II Consider an economy with two consumers, Ben and Joe. There are publP=10-e Ben ? 34 10 Joe 21 P=8-20 P=18-30 Market - MC P=10-e 10ei) How many sirens will be provided in this market? j) What will the price for these sirens be? k) Is the result from previou
Answer #1

Past I

a. If externality is not corrected market equilibrium occurs where Demand is equal to marginal private cost.

100-Q=10+Q

Q=90÷2= 45

Equilibrium Quantity= 45

Equilibrium price= 10+45=55

b. Socially optimal quantity is produced where demand curve meets marginal social cost.

100-Q=10+Q+0.5Q

Q=90/2.5=36

Socially optimal quantity= Q*=36

C. Tax should be equal to the vertical distance between the marginal private cost and marginal social cost.

At Q =36,

Marginal private cost= 10+36= 46

Marginal social cost= 10+1.5*36= 64

Tax= $64-$46= $18 per unit.